It is estimated that 0.42 percent of the callers to the Customer Service department of Dell Inc. will receive a busy signal. What is the probability that of today’s 1,200 callers at least 5 received a busy signal?
Answer by Boreal(15235)

This is a binomial
look at 0,1,2,3,4 and subtract that sum from 1, and the answer will be at least 5.
p(x=0)=9958^1200=0.0064
p(x=1)=0.9958^1199*0.0042*1200=0.0324
p(x=2)=1200C2*0.9958^1198^0.0042^2=0.0820
p(x=3)=1200C3^0.9958^1197*0.0042^3=0.1381
p(x=4)=1200C4*0.9958^1196*0.0042^4=0.1743
sum=0.4332
the answer is therefore 0.5668
Can check on the calculator using 1-binomcdf(1200,0.0042,4) and get 0.5669. There is slight rounding error in the above calculation.